# Bits-to-Photon: End-to-End Learned Scalable Point Cloud Compression for Direct Rendering

### v1

@@ 3.3 From Features to 3D Gaussians
In practice for $N=10$, we found that this generative prediction is helpful at level $N-2$ or lower. But at level $N-1$ and $N$, we just need to predict one Gaussian for each non-empty point while setting the opacity of all Gaussians to a constant of 1. For each Gaussian, we predict the offset from the original grid coordinates and other parameters including $\sigma_{x},\sigma_{y},\sigma_{z},q_{w},q_{x},q_{y},q_{z},c$.
3.4 Training
@@
With the hierarchical coding structure and the scalable delivery and rendering scheme, we encode the point cloud into bit-streams at levels $L$ to $N$, and depending on the target bit rate, deliver from level $L$ up to level $M$ and render at level $M$, with $L\leq M\leq N$. As mentioned earlier, some lower levels are too sparse for adequate rendering, thus they are only used for the conditional coding purpose. Therefore, we have $M>L$ in general.
@@
where the bit-rate term $\mathcal{L}_{R}^{n}$ is defined as in Eq. (5) for level $n$. The distortion loss $\mathcal{L}_{D}^{n}$ includes the weighted L1 distance, the SSIM loss, and the LPIPS loss [32], calculated between the rendered images $\hat{x}^{n}$ at level $n$ and the ground truth views $x$ obtained by rasterizing the mesh provided in the training dataset, defined as
@@
During training, we calculate the loss function using 4 random views rendered from a scene. We set $\alpha=3$, $\beta=0.2$, and $\gamma=1$ in our experiments. Note that to achieve more fine-grained rate control, besides the level scalability, we also train different models with different $\lambda$. We set $\lambda\in\{5,10,20\}$ in our experiments. It thus produces 3 sets of scalable bit-streams. As described earlier, for $N=10$, we choose $M_{\min}=8$ and $M_{\max}=9$, because they lead to better rate-quality trade-offs and adequate rendering quality. Thus each scalable bit stream obtained by a model trained with a given $\lambda$ has 2 rate-quality points, realizing a total of 6 Rate-Quality points. We use the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 4. We train the model for 60,000 iterations on a single NVIDIA A100 GPU.
@@ 4 Experiments
4.1 Settings
We leverage the THuman 2.0 dataset [30] for training and testing, and also evaluate our trained models on the 8i Voxelized Full Bodies (8iVFB) dataset [8]. The THuman 2.0 dataset provides 525 meshes built from RGBD captures with human subjects performing actions. We densely sample the meshes to obtain the color point clouds, and use the point clouds to train and evaluate our model. The dataset is divided into a training set of 500 meshes (00000499($0$-0499( $0$ 0499) and the rest for testing, this guarantees a cross-subject validation with the testing set containing only unseen subjects. In our experiments, we report results with 8 point clouds from unrepeating subjects, namely $\{0507,0509,0511,0513,0515,0517,0519,0521\}$. To evaluate the generalizability, we also test on the MPEG CTC 8iVFB dataset, which provides 4 full-body human point clouds in bit-depth 10, which can be represented losslessly by a 10-level octree.
We evaluate the following compression schemes:
  • G-PCC. We use the MPEG G-PCC reference software TMC13v22 [1] to encode both the geometry (using octree) and the color (using RAHT). It varies the geometry quantization scale and color quantization parameters to achieve different bit-rates. We use the default settings for geometry and color quantization.
  • Learned baseline. We use the state-of-the-art deep learning based scalable point cloud geometry compression method SparsePCGC [25] to encode the geometry, and use the learned point cloud attribute compression method 3DAC [9] to encode the color.
  • B2P (G-PCC). We use our method to compress the point cloud, given geometry coded by G-PCC up to a required level. We use the same settings as the G-PCC method for geometry coding, and use our method for color coding.
  • B2P (SpPCGC). Same as above but we replace the geometry coding with the state-of-the-art deep learning based scalable point cloud geometry compression method SparsePCGC [25].
We report bit-rates in bit-per-point (bpp), using the number of points from the original point cloud. We evaluate the rendering distortion at different bit-rates for comparison. We average the rendering distortion over 12 camera views forming a circle with equal angle intervals around the subject. Our decoder operation includes entropy decoding the squeezed features, reconstructing the features, and generating the Gaussian parameters, which can then be used to efficiently produce the rendered images given the camera extrinsics and intrinsics. For the G-PCC and the learned baseline, we first decode the point cloud, and then render with the existing standard point cloud renderer provided by OpenGL and packaged in Open3D [33], which rasterizes each point as a solid square in the camera plane. We set the square size according to the point density so there is no visible hole. We also include an alternative renderer, taking each point as a spherical Gaussian with a global covariance matrix $\Sigma=\text{diag}(\sigma,\sigma,\sigma)$, where $\sigma=\bar{d}$ corresponding to the average nearest point distance. Since THuman 2.0 dataset provides ground truth meshes, we evaluate the rendering quality using the PSNR, MS-SSIM, and the LPIPS [32] between the rendered images from the decoded point clouds and the same views rendered from the ground truth meshes using standard rasterization. For the 8iVFB dataset which does not provide ground truth meshes, we first use the Screened Poisson surface reconstruction [17] to convert the point clouds to the meshes, and then render the meshes to produce ground truth images.
4.2 Rate-Distortion Performance
@@
Figure 4: Rendering distortion at different bit-rate by the proposed Bits-to-Photon (B2P) and baseline methods. We train 3 models with different $\lambda$ values. Each model has 2 levels of scalability, forming 3 groups of scalable R-D points, plotted using different colors. Other methods are not scalable.
@@
G-PCC (0.62 bpp)
G-PCC (0.62 bpp)
Original
(Groun-Truth Mesh)
B2P (0.49 bpp)
G-PCC (0.53 bpp)
G-PCC (0.53 bpp)
Original
(Poisson Mesh)
B2P (0.59 bpp)
G-PCC (0.62 bpp)
G-PCC (0.62 bpp)
@@
We plot the R-D points achieved by the proposed Bits-to-Photon (B2P) and compared methods in Fig. 4. As shown by the PSNR evaluation, images rendered by our method from our decoded bit-stream have more pixel-level fidelity to the ground truth images rendered from the mesh, with over 4 dB improvements in PSNR at the same level of bit-rate over the compared methods. We also show the results using LPIPS as the distortion metric, which is more sensitive to texture similarity and more tolerant to pixel-wise difference. Despite the disadvantage of Gaussian based renderer over OpenGL in terms of sharpness, our method still achieves better rate-distortion tradeoffs. We note that, from the R-D curves presented, B2P with $\lambda=20$ (orange points) is not an appropriate choice, yielding higher rates at similar quality compared to $\lambda=10$. In Fig. 5 and 6, we demonstrate better visual quality with finer facial details, sharper edges and higher contrast at a lower bit-rate, compared to G-PCC.
4.3 Bit-rate, Complexity, and Distortion Tradeoff
@@
G-PCC 9+ [^2] Using geometry quantization scale $0.75$. Gaussian 0.68 26.5 0.016 0.9854 4 sec 3 ms
@@
Learned 9 OpenGL 0.60 27.7 0.010 0.9893 370 ms 1 sec 3 ms
In a volumetric streaming system, there is usually a tradeoff among the bandwidth consumption, decoding and rendering complexity, and the rendering quality. We study this tradeoff comparing the baseline methods and B2P with the 8iVFB dataset. Our evaluation is conducted on a desktop computer with an Intel i7-9700K CPU and an NVIDIA RTX 4080 Super GPU. B2P, SparsePCGC, Global-Parameter Gaussian-based rendering, 3DAC and Pointersect are using both CPU and GPU, where the neural networks parts are running on GPU. G-PCC is running on CPU.
We show the benchmarking results in Table 1. As shown, B2P shows better R-D performance than compared methods at a lower decoding latency. We have the following observations:
  • B2P decoding and rendering at octree level 8 already achieves better R-D performance than G-PCC at level 9. Although given a better point cloud renderer (Pointersect), the rendering distortion can be further reduced with the same reconstructed point cloud from G-PCC, the rendering complexity is too high for a real-time 6-DoF display.
@@
  • B2P provides better rendering quality by further decode one more level in the octree, as an enhancement layer. Although the decoding time will increase (but still substantially lower than G-PCC and 3DAC), since B2P is scalable, we may conduct region-adaptive enhancement layer decoding in a real application and achieve a better tradeoff between complexity and R-D performance. We leave this as a promising future work.
We further analyze the bit-rate consumption by B2P at each of the hierarchical coding levels (7, 8, and 9), shown in Table 2. We observe that level 8 generally takes the largest portion of the bits, while level 9 does not take as much. This suggests that there are more unpredictable information at level 8 from level 7 features, while level 9 features can be predicted from level 8 features quite well. We will explore improvement of feature prediction from level 7 to level 8 as a future work. On the other hand, the unconditional coding at level 7 consumes about 10% of the bit-rate, but the portion is larger with a lower $\lambda$. This suggests that further improvements can be made to coding the coarser levels to improve performance at a lower bit-rate region.
Table 2: Average bit-rate (bpp) for color feature coding at different levels and $\lambda$.
Dataset $\lambda=5$ $\lambda=10$ $\lambda=20$
Level 7 Level 8 Level 9 Level 7 Level 8 Level 9 Level 7 Level 8 Level 9
THuman 2.0 0.049 0.136 0.010 0.053 0.333 0.092 0.057 0.509 0.182
8iVFB 0.037 0.164 0.007 0.040 0.358 0.086 0.043 0.503 0.175
5 Discussion of Limitations
In this paper, we show that it is promising to jointly design a point cloud compression scheme with a differentiable renderer for lower decoding and rendering latency, and better rate vs.rendering-distortion performance. Here we discuss some limitations of this work and point to possible solutions for future works.
Although our method has certain capability of scalability, it only supports a 2-stage scalable coding and rendering. It may be desirable to achieve a more fine-grained scalability between level 9 and 10 in an actual streaming application with adequate bandwidth. This can be potentially achieved by a region-adaptive coding scheme, which allocates more bits and produces more Gaussians in areas where more points are present. This is also promising for reducing the decoding / rendering latency towards a real-time system.
On the other hand, we observe that the existing 3D Gaussian rasterizer tends to lower the texture contrast (compared to the ground truth) due to the alpha blending mechanism. An earlier work [35] shows that this can be resolved by carefully handling the intersection region of two Gaussians. By improving the differential renderer, we may also be able to boost the R-D performance and visual quality.
@@ 6 Conclusion
Appendix
Appendix 0.A Network Architecture
We provide details on the sparse 3D convolution-based network architecture used in our method. This includes the Feature Convert module, the Feature Squeeze module, the Predictive Entropy Model, the Feature Reconstruct module, and the 3D Gaussian Generation module. Please refer to Fig. 2 (Feature Convert, Feature Squeeze and Predictive Entropy Model) and Fig. 3 (3D Gaussian Generation) in the main paper for how these modules are incorporated in the proposed compression scheme.
0.A.1 Buidling Blocks
(a) InceptionResBlock
(b) ResBlock
Figure 7: The layer-wise details of the InceptionResBlock and ResBlock. SpConv is the abbreviation for 3D sparse convolution (Minkowski Convolution [6]). $C$ is the number of input channels and $N$ is the number of points in the input. The batch dimension is omitted for simplicity.
The proposed architecture includes two kinds of building blocks, i.e. InceptionResBlock and ResBlock. We adopt the InceptionResBlock architecture from the work [27]. For completeness, we also provide the detail in this section. The ResBlock is a 3D sparse version of the original residual block [13]. The layer-wise details of these two blocks are shown in Fig. 7.
0.A.2 Feature Convert Module
Table 3: The architecture of the Feature Convert module.
Layer Layer Type Input Channel Output Channel Kernel Size Activation
1 Geom-Invariant 3D Conv. $C_{\text{in}}$ 64 $3\times 3$ None
2 InceptionResBlock 64 64 - ReLU
3 InceptionResBlock 64 64 - ReLU
4 Geom-Invariant 3D Conv. 64 64 $3\times 3$ None
The Feature Convert module architecture shown in Table 3 is used for both extracting feature from the input point cloud (RGB colors as features) or converting a down-sampled point cloud with features from a denser level to features for this level. The output will 1) go to the Feature Squeeze module for further processing, and 2) be down-sampled to the next level for further processing by another Feature Squeeze module.
0.A.3 Feature Squeeze Module
Table 4: The architecture of the Feature Squeeze module.
Layer Layer Type Input Channel Output Channel Kernel Size Activation
1 Per-point Linear Layer 64 + 64 64 $1\times 1$ ReLU
2 Per-point Linear Layer 64 8 $1\times 1$ None
The Feature Squeeze module architecture shown in Table 4 is used for reducing the dimensionality of the features at this level. Since we want to remove the predictable information to save bit-rates, we design the process to be conditioned on the up-sampled feature from a coarser level. Hence, the input layer in Table 4 has $64+64$ input channels, corresponding to the features at this layer and the context from a coarser layer, respectively. The output will be quantized and entropy coded, with a Predictive Entropy Model.
0.A.4 Predictive Entropy Model
Table 5: The architecture of the Predictive Entropy Model module.
Layer Prev. Layer Type Input Channel Output Channel Kernel Size Activation
1 - Geom-Invariant 3D Conv. 64 64 $3\times 3$ ReLU
2 1 Geom-Invariant InceptionResBlock 64 64 - ReLU
3 2 Geom-Invariant InceptionResBlock 64 64 - ReLU
4 3 Per-point Linear Layer 64 8 $1\times 1$ None
5 3 Per-point Linear Layer 64 8 $1\times 1$ Exp
The Predictive Entropy Model architecture shown in Table 5 is used for modeling the conditional entropy of the quantized features. We assume the squeezed features follow a conditional Gaussian distribution. The Feature Squeeze module thus takes the upsampled feature from a coarser level as input, and generates the mean and scale of the Gaussian distribution for each point, which is used to form the cumulative distribution function (CDF) for entropy coding. In Table 5, Layer 4 generates the mean $\mu$ while Layer 5 generates the scale $\sigma$. The output will be used for entropy coding the quantized features from the Feature Squeeze module.
0.A.5 Feature Reconstruct Module
Table 6: The architecture of the Feature Reconstruct module.
Layer Layer Type Input Channel Output Channel Kernel Size Activation
1 Per-point Linear Layer 8 + 64 64 $1\times 1$ ReLU
2 InceptionResBlock 64 64 - ReLU
3 InceptionResBlock 64 64 - ReLU
4 Per-point Linear Layer 64 64 $1\times 1$ None
The Feature Reconstruct module architecture shown in Table 6 is for reconstructing the higher dimensional feature from the squeezed and quantized features given by the Feature Squeeze module. Since the squeezing removes the predictable information, in the reconstruction the module should take the up-sampled feature from a coarser level as a conditioning input to reconstruct the features. The output will be used 1) as a conditioning input for coding a denser level, and 2) by the 3D Gaussian Generation module to generate the renderable 3D Gaussians.
0.A.6 3D Gaussian Generation
Table 7: The architecture of the 3D Gaussian Generation module.
Layer Layer Type Input Channel Output Channel Kernel Size Activation
1 Geom-Invariant Conv. 64 64 $3\times 3$ ReLU
2 ResBlock 64 64 - ReLU
3 ResBlock 64 64 - ReLU
4 Geom-Invariant Conv. or 3D Transposed Conv. 64 64 $2\times 2$ None
5 ResBlock 64 64 - ReLU
6 ResBlock 64 64 - ReLU
7 Geom-Invariant Conv. 64 14 $3\times 3$ None
The 3D Gaussian Generation module architecture shown in Table 7 is for generating the renderable 3D Gaussians from the reconstructed features. If this module is used at a coarse level (level 8 in our experiments), the 4th layer in Table 7 will be a 3D transposed convolution, which generates 8 times more points than the input for finer details. If this module is used at a denser level (level 9 in our experiments), the 4th layer in Table 7 is configured as a Geometry-Invariant 3D convolution, which generates the same number of points as the input. The output includes 14 channels for the 3D Gaussian parameters, i.e. $\mu_{x},\mu_{y},\mu_{z},\sigma_{x},\sigma_{y},\sigma_{z},q_{w},q_{x},q_{y},q_{z},o,c$.
Appendix 0.B Discussion on Point Cloud Video Compression and Rendering
In this section, we discuss a promising extension of our method to point cloud video compression and rendering. We conduct this experiment with the 8iVFB dataset [8]. We compare our method to the standard scheme, whice employs G-PCC for point cloud compression and OpenGL for rendering. Please kindly refer to the supplementary video for a demonstration of results.
As shown, our method is able to achieve a better quality at a lower bit-rate compared to the standard scheme. More importantly, when we look closer to the 3D surface, the holes / gaps between points will be visible in the standard scheme, which degrade the visual quality. In contrast, our method produces 3D Gaussians that approximate the smooth surface. Therefore, it does not suffer from the same issue and shows less visual artifacts.
Nevertheless, since the main focus of this paper is on joint point cloud compression and rendering for one frame, our method does have limitations in point cloud video compression and rendering. For example, we didn’t consider the temporal coherence between frames, which is important for both compression and reconstruction. We notice some flickering effects in the rendered video, which is caused by the lack of temporal coherence. We believe that our method can be extended to point cloud video compression and rendering by incorporating temporal consistency loss and temporal modeling designs. For example, we can use warped previous frames as a conditioning input for the Feature Squeeze and Feature Reconstruct modules to reduce the bit-rates. We can also use this information to improve the temporal coherence in the rendering process, forming spatio-temporal 4D Gaussians. We leave these extensions as a promising future work.

### v2

@@ 3 Method
In practice for $N=10$, we found that this generative prediction is helpful at level $M=N-2$ or lower. But at level $N-1$ and $N$, we just need to predict one Gaussian (i.e. $k=1$) for each occupied point while setting the opacity of all Gaussians to a constant of 1. For each Gaussian, we predict the offset from the original grid coordinates and other parameters including $\sigma_{x},\sigma_{y},\sigma_{z},q_{w},q_{x},q_{y},q_{z},c$.
Training
@@
With the hierarchical coding structure and the scalable delivery and rendering scheme, we encode the point cloud into bit-streams at levels $L$ to $N$, and depending on the target bit rate, deliver from level $L$ up to level $M$ and render at level $M$, with $L\leq M\leq N$.
@@
where the bit-rate term $\mathcal{L}_{R}^{n}$ is defined as in Eq. (5) for level $n$. The distortion loss $\mathcal{L}_{D}^{n}$ includes the weighted L1 distance, the SSIM loss, and the LPIPS loss (Zhang et al. 2018), calculated between the rendered images $\hat{x}^{n}$ at level $n$ and the ground truth views $x$ obtained by rasterizing the mesh provided in the training dataset, defined as
@@
During training, we calculate the rendering distortion averaged from 4 random view points for each point cloud in the training set. We set $\alpha=3$, $\beta=0.2$, and $\gamma=1$ in our experiments. Note that to achieve more fine-grained rate control, besides the level scalability, we also train different models with different $\lambda$. We set $\lambda\in\{5,10\}$ in our experiments. It thus produces 2 sets of scalable bit-streams.
We have observed that, when the level of details is too sparse (i.e, $M$ is too small), the features do not carry sufficient information for rendering with adequate quality. We found that for a typical 10-bit point cloud ($N=10$), the best rate-quality trade-off is achieved by transmitting starting from level $L=7$ up to level $M$ with $M_{\min}=8,M_{\max}=9$. Although it is possible to transmit all way to level $10$, the quality improvement from level $9$ to level $10$ does not justify the additional bits needed. Thus each scalable bit stream obtained by a model trained with a given $\lambda$ has 2 rate-quality points (corresponding to $M=8,9$), realizing a total of 4 Rate-Quality points. When decoding at $M=8$, we use transpose convolution to generate 8 Gaussians for each point. But at $M=9$, we only generate one Gaussian per point. We use the Adam optimizer with a learning rate of $10^{-4}$ and a batch size of 4. We train the model for 60,000 iterations on a single NVIDIA A100 GPU. We will release our code and the trained model to the public upon acceptance.
@@ 4 Experiments
Figure 3: Rendering distortion at different bit-rate by the proposed Bits-to-Photon (B2P) and baseline methods. We train 2 models with 2 different $\lambda$ values. Each model has 2 levels of scalability, forming 2 groups of scalable R-D points, plotted using different colors. Other methods are not scalable.
Settings
We leverage the THuman 2.0 dataset (Yu et al. 2021) for training and testing, and also evaluate our trained models on the 8i Voxelized Full Bodies (8iVFB) dataset (d’Eon et al. 2017). The THuman 2.0 dataset provides 525 meshes built from RGBD captures with human subjects performing actions. We densely sample the meshes to obtain the color point clouds, and use the point clouds to train and evaluate our model. The dataset is divided into a training set of 500 meshes (00000499($0$-0499( $0$ 0499) and the rest for testing, this guarantees a cross-subject validation with the testing set containing only unseen subjects. In our experiments, we report results with 8 point clouds from unrepeating subjects, namely $\{0507,0509,0511,0513,0515,0517,0519,0521\}$. To evaluate the generalizability, we also test on the MPEG CTC 8iVFB dataset, which provides 4 full-body human point clouds in bit-depth 10, which can be represented losslessly by a 10-level octree.
We evaluate the following compression schemes:
  • G-PCC. We use the MPEG G-PCC reference software TMC13v22 (tmc 2023) to encode both the geometry (using octree) and the color (using RAHT). It varies the geometry quantization scale and color quantization parameters to achieve different bit-rates. Therefore the resulting bit points are not part of a scalable stream. We use the default settings for geometry and color quantization.
  • 3DAC. We use the state-of-the-art learned scalable point cloud geometry compression method SparsePCGC (Wang et al. 2022) to encode the geometry, and use the learned point cloud attribute compression method 3DAC (Fang et al. 2022) to encode the color.
  • SAC. Similarly we use SparsePCGC to encode the geometry and use the learned scalable point cloud attribute compression (SAC) model (Zhang et al. 2023b) to encode the color. [^2] Since the code is not publicly available, we use the bit-stream and reconstructions kindly shared by the authors, which only include point clouds in the 8iVFB dataset with lossless geometry.
  • B2P (G-PCC). We use our method to compress the color and other rendering information, given geometry coded by G-PCC up to the same octree level as the color bit-stream. required level. We use the same settings as the G-PCC method for geometry coding.
  • B2P (SpPCGC). Same as above but we replace the geometry coding with SparsePCGC (Wang et al. 2022).
B2P
(0.93 bpp)
G-PCC (1.00 bpp)
3DAC (1.18 bpp)
Original
(GT Mesh)
We report bit-rates in bit-per-point (bpp), using the number of points from the original point cloud. We evaluate the rendering distortion at different bit-rates for comparison. We average the rendering distortion over 12 camera views forming a circle with equal angle intervals around the subject. Our decoder operation includes entropy decoding the squeezed features, reconstructing the features, and generating the Gaussian parameters, which can then be used to efficiently produce the rendered images given the camera extrinsics and intrinsics. For the G-PCC and the learned baseline, we first decode the point cloud, and then render with the existing standard point cloud renderer provided by OpenGL and packaged in Open3D (Zhou, Park, and Koltun 2018), which rasterizes each point as a solid square in the camera plane. We set the square size according to the average point density so there is no visible hole. We also include an alternative renderer, taking each point as a spherical Gaussian with a fixed diagonal covariance matrix $\Sigma=\text{diag}(\sigma,\sigma,\sigma)$, where $\sigma=\bar{d}$ corresponding to the average nearest point distance. Since THuman 2.0 dataset provides ground truth meshes, we evaluate the rendering quality using the PSNR, MS-SSIM, and the LPIPS (Zhang et al. 2018) between the rendered images from the decoded point clouds and the same views rendered from the ground truth meshes using standard rasterization. For the 8iVFB dataset which does not provide ground truth meshes, we first use the Screened Poisson surface reconstruction (Kazhdan and Hoppe 2013) to convert the point clouds to the meshes, and then render the meshes to produce ground truth images.
Rate-Distortion Performance
@@
SAC (0.64 bpp)
@@
G-PCC 9+ [^3] Using geometry quantization scale $0.75$. Global Gaussian 0.68 26.5 0.016 0.9854 4 sec 3 ms
@@
We plot the R-D points achieved by the proposed Bits-to-Photon (B2P) and compared methods in Fig. 3. As shown by the PSNR evaluation, images rendered by our method from our decoded bit-stream have more pixel-level fidelity to the ground truth images rendered from the mesh, with over 4 dB improvements in PSNR at the same level of bit-rate over the compared methods. We also show the results using LPIPS (Zhang et al. 2018) as the distortion metric, which is more sensitive to texture similarity and more tolerant to pixel-wise difference. Despite the disadvantage of Gaussian based renderer over OpenGL in terms of sharpness, our method still achieves better rate-distortion tradeoffs. In Fig. 4 and 5, we demonstrate better visual quality with finer facial details, sharper edges and higher contrast at a lower bit-rate, compared to G-PCC, 3DAC, an SAC.
Bit-rate, Complexity, and Distortion Tradeoff
Table 1 compares rate, rendering quality and complexity (in terms of running time) of different methods with the 8iVFB dataset. Our evaluation is conducted on a desktop computer with an Intel i7-9700K CPU and an NVIDIA RTX 4080 Super GPU. B2P, SparsePCGC, Global-Parameter Gaussian-based rendering, 3DAC and Pointersect are using both CPU and GPU, where the neural networks parts are running on GPU. G-PCC is running on CPU. As shown, B2P shows better R-D performance than compared methods at a lower decoding latency. We have the following observations:
  • B2P decoding and rendering at octree level 8 already achieves better R-D performance than G-PCC at level 9. Although given a better point cloud renderer (Pointersect), the rendering distortion can be further reduced with the same reconstructed point cloud, the rendering complexity is too high for real-time 6-DoF viewing.
@@
  • B2P provides better rendering quality by further decode one more level in the octree, as an enhancement layer.
Ablation Study
[…]
Figure 6: Ablation study on the conditional feature squeeze and reconstruction modules. With B2P (Unconditional), we remove the conditioning in both the feature squeezing and reconstruction module from the coarser layer.
Unlike existing work of learned point cloud attribute compression that either rely on G-PCC to compress a base layer point cloud attributes (Zhang et al. 2023b, a) or only has conditional entropy coding (Wang and Ma 2022), we adopt a hierarchical compression scheme where not only the entropy coding module but also the feature squeezing and reconstruction modules are conditioned on a coarser layer that has been already coded. We conduct an ablation study to show the efficacy of this technique by comparing the full scheme to a version disabling the conditional feature squeezing and reconstruction. As shown in Fig. 6, the conditional feature squeezing and reconstruction modules greatly contributed to the rate-distortion performance by reducing more than 25% of the bit-rate at the same level of reconstruction quality. This technique can be also adopted by other learned scalable point cloud compression schemes for potential performance boost.
@@ 5 Conclusion
We have only demonstrated promising results with 2 scalable levels so far even though the framework is capable of supporting more levels. It may be desirable to achieve a more fine-grained scalability between level 9 and 10. This can be potentially achieved by a region-adaptive coding scheme, which allocates more bits and produces more Gaussians in areas where more points are present at level 10. Furthermore, we have only considered coding a single point cloud. Future research will consider coding point cloud video by exploiting decoded features in previous frames as temporal context for feature squeezing, entropy coding and reconstruction.
Acknowledgement
This material is based upon work supported by the National Science Foundation under Grant No. 2312839.